OpenStudy (anonymous):

The formula for degrees Celsius is: C = 9 over 5 (F – 32), where F stands for degrees Fahrenheit. Part 1: Solve the formula for F. Show all steps. Part 2: Determine how many degrees Fahrenheit 6 degrees Celsius is

OpenStudy (anonymous):

Isolate F the same way C is isolated above. \[C = \frac{ 9 }{ 5 } (F - 32)\]\[\frac{ 5 }{ 9 }C = F- 32\]\[F = 32 + \frac{ 5 }{9 }C\] for part 2, plug C = 6 to find F